what is the mass of oxygen molecule where molecular mass is 32?
1
Expert's answer
2014-10-28T11:05:43-0400
The mass of the molecule is:
m(O2) = 32*1.660538921(73)×10−27 = 5.31×10−26 kg = 5.31×10−23 g